When Frank and Mark finished milking, they found that they had 900 pounds of milk that was 2 percent butterfat. How much butterfat did they have to add to raise the butterfat content to 10 percent?
Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

X+.02*900=.10(900+X)
X+18=90+.10X
.9X=90-18
.9X=72
X=72/.9
X=80 GALLONS OF BUTTERMILK IS NEEDED.

Answer by stanbon(48558)

When Frank and Mark finished milking, they found that they had 900 pounds of milk that was 2 percent butterfat. How much butterfat did they have to add to raise the butterfat content to 10 percent?
-----------
Amt. of 2% milk is 900 lbs; amt. of fat is 0.02(900)=18 lbs
------------
Amt of 100% butterfat mild is x: amt. of fat is x lbs
--------
Amt of 10% mixture is (900+x); amt of fat is 0.10(900+x)=90+0.1x lbs
----------
EQUATION:
fat + fat = fat
18 + x = 90 + 0.1x
0.9x = 72
x = 80 lbs (amount of 100% butterfat that must be added)
-----------
Cheers,
Stan H.